OpenStudy (anonymous):

The diameter of the larger circle is 12.5 cm The diameter of the smaller circle is 3.5 cm What is the approximate area of the shaded region?

OpenStudy (anonymous):

Area of larger circle-area of smaller circle

OpenStudy (anonymous):

Area of large circle:: \[\pi *r _{1}^{2}\] where r1 is the radius of lager circle obtained by halving the diameter

OpenStudy (anonymous):

it does not give me that...?

OpenStudy (anonymous):

can u please give me the answer pleaseee

OpenStudy (anonymous):

{(3.14*12.5*12.5)/4} - {((22/7) *(3.5)*(3.5))/4}

OpenStudy (anonymous):

how would i do that with a fraction..?

OpenStudy (anonymous):

Simply solve it....

OpenStudy (anonymous):

254.35

OpenStudy (anonymous):

(490.625)/4 - (38.465)/4

OpenStudy (anonymous):

113.04

OpenStudy (anonymous):

132.27

OpenStudy (anonymous):

one more...?

OpenStudy (anonymous):

another question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ya tell me

OpenStudy (anonymous):

What is the approximate area of the shaded region? 

OpenStudy (anonymous):

the square is the shaded area

OpenStudy (anonymous):

Area of the square=(Side)^2

OpenStudy (anonymous):

7cm right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Area of circle : \[\pi*r*r \] where r=3.5cm

OpenStudy (anonymous):

then subtract

OpenStudy (anonymous):

38.465...? right

OpenStudy (anonymous):

subtract what...?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

area of square - area of circle

OpenStudy (anonymous):

can u give me the numbers pleasee

OpenStudy (anonymous):

49- 38.465

OpenStudy (anonymous):

10.535

OpenStudy (anonymous):

what next

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no issues....